OpenStudy (anonymous):

is it correct? 4x^2-y^2+2y-1 = 4x^2-(y-1)^2 =(2x-y-1)(2x+y-1) 16p^2-(5q-r)^2 = 16p^2=(4p)(4p) 5q-r^2=(5q-r)(5q+r) =(4p+5q-r)(4p-5q+r)

OpenStudy (perl):

almost

OpenStudy (perl):

4x^2-y^2+2y-1 = 4x^2-(y-1)^2 = ( 2x - (y-1) ) ( 2x + (y-1) ) = ( 2x - y + 1 ) ( 2x + y - 1)

OpenStudy (anonymous):

the other one is correct?

OpenStudy (perl):

\[16p^2-(5q-r)^2 = (4p - (5q-r)) (4p + (5q-r)) = ( 4p - 5q + r) (4p +5q -r) \]

OpenStudy (perl):

im sorry you did it right :D

OpenStudy (anonymous):

thank you perl im just making sure, im not confident with my answer and sometimes i forget how to do it. specially in factoring and groupings.

OpenStudy (perl):

this is not right however 5q-r^2=(5q-r)(5q+r) !!! false

OpenStudy (anonymous):

ok then ill try to rewrite it.

OpenStudy (perl):

oh i see the problem

OpenStudy (perl):

it said ( 5q - r)^2 , not 5q-r^2 , which is different

OpenStudy (perl):

\[16p^2-(5q-r)^2 \\= (4p - (5q-r)) (4p + (5q-r)) \\= ( 4p - 5q + r) (4p +5q -r)\]

OpenStudy (anonymous):

thank you perl i got it, but im confused to this 2y^2-3z = what supposed to be the answer?

OpenStudy (perl):

you should just leave it be :)

OpenStudy (perl):

\[2y^2-3z \]

OpenStudy (anonymous):

cant it be (y+3z)(y-3z)?